# Dual Graph Attention based Disentanglement Multiple Instance Learning for Brain Age Estimation

### v1

@@ Abstract
Deep learning techniques have demonstrated great potential for accurately estimating brain age by analyzing Magnetic Resonance Imaging (MRI) data from healthy individuals. However, current methods for brain age estimation often directly utilize whole input images, overlooking two important considerations: 1) the heterogeneous nature of brain aging, where different brain regions may degenerate at different rates, and 2) the existence of age-independent redundancies in brain structure. To overcome these limitations, we propose a Dual Graph Attention based Disentanglement Multi-instance Learning (DGA-DMIL) framework for improving brain age estimation. Specifically, the 3D MRI data, treated as a bag of instances, is fed into a 2D convolutional neural network backbone, to capture the unique aging patterns in MRI. A dual graph attention aggregator is then proposed to learn the backbone features by exploiting the intra- and inter-instance relationships. Furthermore, a disentanglement branch is introduced to separate age-related features from age-independent structural representations to ameliorate the interference of redundant information on age prediction. To verify the effectiveness of the proposed framework, we evaluate it on two datasets, UK Biobank and ADNI, containing a total of 35,388 healthy individuals. Our proposed model demonstrates exceptional accuracy in estimating brain age, achieving a remarkable mean absolute error of 2.12 years in the UK Biobank. The results establish our approach as state-of-the-art compared to other competing brain age estimation models. In addition, the instance contribution scores identify the varied importance of brain areas for aging prediction, which provides deeper insights into the understanding of brain aging.
@@ 1 Introduction
Aging is commonly defined in most studies as a temporal decline in function that affects a wide range of organisms López-Otín et al. (2013); Adler et al. (2007); Cole et al. (2019). While great efforts have been made towards understanding aging, the complex biology of the aging process remains to be fully characterised López-Otín et al. (2013). As individuals age, cognitive decline becomes more prevalent, increasing the risk of neurodegenerative disorders and dementia. However, the aging process itself is hardly observable directly. Instead, it can often be inferred from quantifiable accompanying phenomena. In the field of biogerontology, researchers have long been searching for these quantifiable features, known as biomarkers of aging Cole et al. (2019). It’s well recognized that throughout the aging process, there are notable changes in both the morphology and function of tissues and organs within the body Cole et al. (2019). Specifically, in the brain, aging induces several physiological changes such as ventricular enlargement, cortical thinning, and the accumulation of white matter hyperintensities Terry et al. (1987); Fjell et al. (2014); Fotenos et al. (2005). Consequently, there is an opportunity to use neuroimaging technologies for measuring brain age and capturing structural changes in the brain.
@@
Deep learning techniques have achieved remarkable success in the field of medical image analysis, primarily due to their proficiency in learning comprehensive features from clinical imaging datasets. These methods have been applied in a wide range of areas, including object segmentation, disease diagnosis, and brain age estimation He et al. (2021b); Chen et al. (2023). There has been an increasing interest on estimating brain age using deep learning approaches on structural magnetic resonance imaging (sMRI) Liu et al. (2017); Madan and Kensinger (2016); Liem et al. (2017); Lin et al. (2016); Xu et al. (2021); Cherubini et al. (2016); Feng et al. (2020); Bashyam et al. (2020); Dinsdale et al. (2021); He et al. (2021b); Lam et al. (2020); He et al. (2022); Beheshti et al. (2021); Ganaie et al. (2022). sMRI is capable of capturing changes in brain anatomy and morphological atrophy. Compared to other biomedical imaging techniques, sMRI allows non-invasive examination of a wide range of structural and functional aspects of the human brain, making it a widely utilized tool in both brain aging research and clinical studies of brain disorders Cherubini et al. (2016); Feng et al. (2020); Bashyam et al. (2020); Dinsdale et al. (2021); He et al. (2021b). Typically, individuals without neurological disorders are assumed to have a normal brain aging trajectory, resulting in an approximate congruence between their brain age and chronological age. To achieve accurate prediction of brain age, it is common practice to train a deep learning model with sMRI data from healthy individuals, using their chronological age as the target.
@@
Recent advances in brain age estimation have seen the successful exploration of deep learning techniques in MRI analysis. For example, BrainNet Bashyam et al. (2020) independently predicts an individual’s age using 2D slices derived from 3D sMRI, without considering the relationships between slices on age prediction. Recurrent CNN Lam et al. (2020) takes the extracted slice features by a convolutional neural network (CNN) as input to a recurrent neural network. It learns the inter-slice information but fails to provide interpretability of each slice’s contribution to the age prediction. In another line of study, Graph Transformer Cai et al. (2022) is proposed, which constructs brain networks to learn regional geometric information and obtain age-related features specifically within local regions of interest (ROI) for age prediction. However, it does not take into account the global information in non-ROI regions, thereby possibly compromising the accuracy of the overall prediction. In addition, a Simple Fully Convolutional Network (SFCN) Peng et al. (2021) employs a 3D CNN framework to capture whole image features for age prediction and yields improved estimation results.
@@
While deep learning-based models have greatly enhanced the performance of brain age estimation, the models using the entire 3D MRI as input may ignore the inhomogeneity in aging across different brain areas, making the network focus solely on salient information and thus potentially ignoring finer details Cole and Franke (2017). Recent evidence suggests that age-related metabolism is most pronounced in the frontal lobes, posterior cingulate, and posterior parietal lobes, while the thalamus, putamen, and pallidum are less susceptible to age-related changes Lee et al. (2022); Kaufmann et al. (2019); Zhu et al. (2023). In addition, Man et al. report that brain age is associated with brain volume, including subcortical regions and the prefrontal cortex Zhu et al. (2023). These prior studies indicate that different brain regions may undergo varying degrees of degeneration. In other words, some regions exhibit significant age-related changes, while others show relatively minor. On the contrary, a number of models rely on 2D slices as input. However, they treat each slice independently and average the model outputs for prediction which overlooks the inter-slice relationships and often yields suboptimal results. Therefore, it is critical to build a model for capturing the unique patterns of aging across the brain with sufficient consideration of inter- and intra-slice features. Additionally, sMRI can contain a wealth of anatomical and morphological information, while only a portion of them is directly relevant to aging. The redundant features may interfere with the estimation of age, leading to degraded performance.
@@
Based on the above observations, we propose a novel end-to-end architecture called Dual Graph Attention based Disentanglement Multiple Instance Learning (DGA-DMIL), specifically designed for accurate brain age estimation using sMRI data. It takes multiple 2D slices in sMRI as input and uniformly groups them as a bag, which is fed into a 2D CNN backbone module to capture spatial detail within each instance. A dual graph attention aggregator is then proposed to combine the spatial backbone features of inter- and intra-instances. The dual graph attention aggregator is capable of assigning different weights to different instances, thereby addressing the issue of brain heterogeneity in aging. Furthermore, we introduce a disentanglement branch to separate age-related features from age-independent structural representations to prevent the negative impact of redundant information on age prediction.
The effectiveness of the proposed model is extensively validated on two datasets, UK Biobank and ADNI, with a total of 35,388 samples which is one of the largest sample sizes for brain age estimation. It achieves a mean absolute error (MAE) of 2.12 years in UK Biobank and 2.81 years in ADNI, surpassing the performance of existing state-of-the-art approaches. In addition, our method proves valuable in identifying slices that significantly contribute to brain age estimation, facilitating the localization of the age-related regions. In contrast to prior studies, the main contributions of this work can be summarised as follows:
@@
    We propose a novel end-to-end multi-instance learning framework, called DGA-DMIL, to fully exploit intra- and inter-instance relationships for accurate brain age estimation.
@@
    To address the challenge of heterogeneous aging, we introduce a dual graph attention aggregator to evaluate the importance of different instances of aging and to aggregate instance features for improved brain age estimation. In addition, we use instance contribution scores to visualize the impact of different regions on aging.
@@
    To ameliorate the influence of redundant information, we propose a disentanglement branch that separates age-related and age-irrelevant features to prevent the adverse effect of redundant information on age prediction.
@@ 2.1 Brain age estimation
With superior feature representation abilities, deep learning-based techniques have been successfully applied to medical image analysis in various medical fields including brain age estimation Cherubini et al. (2016); Feng et al. (2020); Bashyam et al. (2020); Dinsdale et al. (2021); He et al. (2021b); Kassani et al. (2019); Hu et al. (2019). Most existing brain age estimation methods rely on CNN or Transformer architectures to extract features from 2D MRI slices or 3D volumes. One notable approach adopts the two-stage age network (TSAN) to estimate a rough age at the first stage, followed by a second stage that refines the estimation for improved accuracy Cheng et al. (2021). This approach achieves an MAE of 2.43 years on a dataset of 6,586 samples. Another lightweight network, SFCN, proposes an architecture without fully connected layers in brain estimation Peng et al. (2021) and achieves an MAE of 2.14 years on a larger dataset of 14,503 samples. In addition, a global-local transformer model is introduced to capture both global contexts from the entire slices and local fine-grained features from specific patches He et al. (2021a). By using attention mechanisms to combine these features, it achieves an MAE of 2.70 years on a dataset of 8,379 samples.
Notably, in this study, our proposed method allows us to exploit the attention scores within and between multiple brain slices to estimate brain age. In addition, it introduces a decoupling method that effectively separates age-related features from age-irrelevant structural features, further improving the performance of age prediction. Our method achieved an MAE of 2.12 years on the UK Biobank dataset of 35,291 individuals, outperforming existing approaches.
@@ 2.2 Multi-instance learning
Multi-instance learning (MIL) is a type of weakly supervised learning in which labels are only assigned to a bag of instances, while the labels of instances within the bag remain unknown Zhou (2018). MIL can accurately predict the label of the bag by taking into account the effect of the instance on the final result and provide insights on the impacts of intra- and inter-slice features Han et al. (2020); Chikontwe et al. . Thus, it has been widely applied in various domains, including medical imaging Shao et al. (2021). Ilse et al. propose an attention-based MIL pooling mechanism focusing on cancer detection in histopathological images Ilse et al. (2018). Marx et al. apply MIL to predict age based on the hippocampal slices Marx et al. (2023). However, it is limited to the hippocampal region, neglecting important information in other brain areas. To the best of our knowledge, our approach is the first attempt to use the MIL method for age prediction based on brain sMRI data. Unlike other MIL models, we perform dual graph attention-based aggregation on the deep features extracted from each instance and on multiple instances from the bag. We accurately predict the age of the brain by aggregating the information obtained within and across instances.
@@ 2.3 Decoupled Representation Learning
Disentangled representation learning aims to capture the explanatory factors from diverse data variations, and has attracted considerable attention in recent years. It has shown great promise in the field of computer vision Kinney and Atwal (2014); Hou et al. (2021). Previous studies have used annotated data to separate representations into task-relevant and task-independent information Hou et al. (2021). One notable approach, InfoGAN Chen et al. (2016), maximizes the mutual information between hidden and data variables to successfully disentangle latent factors and thus improve performance. Recent work by Hu et al. Hu et al. (2020) introduced the Adversarial Autoencoder framework to decouple shared and complementary features across different modalities, intending to accurately predict brain age by providing specific explanations of the decomposition factors. There are notable differences in our proposed method with Hu et al. (2020). (1) Architecture: We adopt an end-to-end non-generative network framework. (2) Decoupling goal: Our research focuses on decoupling age information and non-age information, which is particularly relevant for the downstream task.
@@ 3.1.1 Problem Formulation
In traditional classification or regression tasks, the goal is to build a model that can accurately predict the target variable, denoted as $y\in R$, given an instance represented by $x\in R^{D}$ Ilse et al. (2018). However, in the context of MIL, the input to the model is a collection of instances called a bag, which is represented as $X_{i}=\left\{x_{i,1},x_{i,2},...,x_{i,k}\right\}\in R^{K*D}$, where K is the number of instances within the bag, and D represents the dimension of each instance. There exists a label $y\in R$ at the bag level. The main goal is to gain insight into the inherent properties and contributions of individual instances within a collection of instances for the overall label assigned to the collection. Ultimately, the model aims to make predictions about the label $y$ for the given collection of instances. The formulation of MIL for a regression task is expressed as,
@@ 3.2 CNN backbone
To analyze sMRI data, we adopt a CNN network as the underlying architecture to extract initial features from instances, as shown in Figure 2. To prevent potential overfitting issues associated with deep networks, we employ the VGG13Simonyan and Zisserman (2014) model as our CNN backbone, which is a well-established architecture widely used in medical imaging. The VGG13 model consists of eight blocks, each comprising a $3\times 3$ convolutional layer with 1 padding, a batch normalization layer, and a ReLU activation layer. After every two blocks, a maximum pooling layer with a $2\times 2$ kernel size and a stride of 2 is applied to gradually reduce the size of the spatial feature map. The number of channels utilized in each block is $[64,128,256,512]$.
@@ 3.2.2 Dual Graph Attention Aggregator
Inspired by the attention-based MIL pooling method developed in Ilse et al. (2018), we propose a novel MIL module, the dual graph attention aggregator (as illustrated in Figure 4), building on the graph attention mechanism. The dual graph attention aggregator consists of a spacial-level feature aggregation network $A_{\theta,S}(R^{K*C*H*W}\to R^{K*C})$ and an instance-level feature aggregation network $A_{\theta,I}(R^{K*C}\to R^{C})$. In the aggregator $A_{\theta,S}$, our goal is to aggregate spatial features from the feature map $e_{i,j}$ into channel features. By passing the feature map $e_{i,j}$ through the GAT block, which consists of GAT, LN, and full connection (FC) layers to reduce the channel dimension to 1, and a softmax layer for normalization, we obtain the final spatial score $g_{i,j}$. To compute the spatial score $g_{i,j}\in R^{1*H*W}$, we perform an element-wise multiplicative convolution of each branch, taking the output layers of the previous layer as input. Mathematically, $g_{i,j}^{{}^{\prime}}=g_{i,j}*e_{i,j}\in R^{C}$, which we denote as $g_{i,j}$ for consistency. In particular, our approach differs from simple global average pooling of the initial features as we adopt the graph attention mechanism to assign different weights within the spatial feature map.

### v2

# Dual Graph Attention-based Disentanglement Multiple Instance Learning for Brain Age Estimation

@@ Abstract
Deep learning techniques have demonstrated great potential for accurately estimating brain age by analyzing Magnetic Resonance Imaging (MRI) data from healthy individuals. However, current methods for brain age estimation often directly utilize whole input images, overlooking two important considerations: 1) the heterogeneous nature of brain aging, where different regions contribute to the aging process at different degrees, and 2) the existence of age-independent redundancies in brain structure. To overcome these limitations, we propose a Dual Graph Attention-based Disentanglement Multi-instance Learning (DGA-DMIL) framework for improving brain age estimation. Specifically, the 3D MRI data, treated as a bag of instances, is fed into a 2D convolutional neural network backbone, to capture the aging patterns in MRI. A dual graph attention aggregator is then proposed to learn the backbone features by exploiting the intra- and inter-instance relationships to provide the instance contribution scores explicitly. Furthermore, a disentanglement branch is introduced to separate age-related features from age-independent structural representations to ameliorate the interference of redundant information on age prediction. To verify the effectiveness of the proposed framework, we evaluate it on two datasets, UK Biobank and ADNI, containing a total of 35,388 healthy individuals. Our proposed model demonstrates exceptional accuracy in estimating brain age, achieving a remarkable mean absolute error of 2.12 years in the UK Biobank. The results establish our approach as state-of-the-art compared to other competing brain age estimation models. In addition, the instance contribution scores identify the varied importance of brain areas for aging prediction, which provides deeper insights into the understanding of brain aging.
@@ 1 Introduction
Aging is commonly defined in most studies as a temporal decline in function that affects a wide range of organisms [38, 1, 12]. While great efforts have been made towards understanding aging, the complex biology of the aging process remains to be fully characterised [38]. As individuals age, cognitive decline becomes more prevalent, increasing the risk of neurodegenerative disorders and dementia. However, the aging process itself is hardly observable directly. Instead, it can often be inferred from quantifiable accompanying phenomena. In the field of biogerontology, researchers have long been searching for these quantifiable features, known as biomarkers of aging [12]. It’s well recognized that throughout the aging process, there are notable changes in both the morphology and function of tissues and organs within the body [12]. Specifically, in the brain, aging induces several physiological changes such as ventricular enlargement, cortical thinning, and the accumulation of white matter hyperintensities [46, 15, 16]. Consequently, there is an opportunity to use neuroimaging technologies for measuring brain age and capturing structural changes in the brain.
@@
Deep learning techniques have achieved remarkable success in medical image analysis, primarily due to their proficiency in learning comprehensive features from clinical imaging datasets. These methods have been applied in a wide range of areas, including object segmentation, disease diagnosis, and brain age estimation [22, 6]. There has been an increasing interest on estimating brain age using deep learning approaches on structural magnetic resonance imaging (sMRI) [37, 39, 35, 36, 48, 9, 14, 3, 13, 22, 33, 20, 4, 17]. sMRI is capable of capturing changes in brain anatomy and morphological atrophy. Compared to other biomedical imaging techniques, sMRI allows non-invasive examination of a wide range of structural and functional aspects of the human brain, making it a widely utilized tool in both brain aging research and clinical studies of brain disorders [9, 14, 3, 13, 22]. Typically, individuals without neurological disorders are assumed to have a normal brain aging trajectory, resulting in an approximate congruence between their brain age and chronological age. To achieve accurate prediction of brain age, it is common practice to train a deep learning model with sMRI data from healthy individuals, using their chronological age as the target.
@@
Recent advances in brain age estimation have seen the successful exploration of deep learning techniques in MRI analysis. For example, BrainNet [3] independently predicts an individual’s age using 2D slices derived from 3D sMRI, without considering the relationships between slices on age prediction. Recurrent CNN [33] takes the extracted slice features by a convolutional neural network (CNN) as input to a recurrent neural network. It learns the inter-slice information but fails to provide interpretability of each slice’s contribution to the age prediction. In another line of study, Graph Transformer [5] is proposed, which constructs brain networks to learn regional geometric information and obtain age-related features specifically within local regions of interest (ROI) for age prediction. However, it does not take into account the global information in non-ROI regions, thereby possibly compromising the accuracy of the overall prediction. In addition, a Simple Fully Convolutional Network (SFCN) [41] employs a 3D CNN framework to capture whole image features for age prediction and yields improved estimation results.
@@
While deep learning-based models have greatly enhanced the performance of brain age estimation, many models using the entire 3D MRI as input may ignore the inhomogeneity in aging across different brain areas, making the network focus solely on salient information and thus potentially ignoring finer details [11]. Recent evidence suggests that age-related metabolism is most pronounced in the frontal lobes, posterior cingulate, and posterior parietal lobes, while the thalamus, putamen, and pallidum are less susceptible to age-related changes [34, 30, 51]. In addition, Man et al. report that brain age is associated with brain volume, including subcortical regions and the prefrontal cortex [51]. These prior studies indicate that some regions exhibit significant age-related changes, while others show relatively minor. In other words, different regions contribute to the aging process at different degrees. On the contrary, many models rely on 2D slices as input. However, they treat each slice independently and average the model outputs for prediction which overlooks the inter-slice relationships and often yields suboptimal results. Therefore, it is critical to build a model for capturing the unique patterns of aging across the brain with sufficient consideration of inter- and intra-slice features. Additionally, sMRI can contain a wealth of anatomical and morphological information, while only a portion of them is directly relevant to aging. The redundant information may interfere with the estimation of age, leading to degraded performance.
@@
Based on the above observations, we propose a novel end-to-end architecture called Dual Graph Attention-based Disentanglement Multiple Instance Learning (DGA-DMIL) specifically designed for accurate brain age estimation. It takes multiple 2D slices in sMRI as input and uniformly groups them as a bag of instances, which is fed into a 2D CNN backbone module to capture spatial detail within each instance. A dual graph attention aggregator is then proposed to combine the spatial backbone features of inter- and intra-instances. The aggregator is capable of weighing different contributions of instances to the brain aging process, thereby addressing the issue of brain heterogeneity in aging. Furthermore, we introduce a disentanglement branch to separate age-related features from age-independent features to prevent the negative impact of redundant information on age prediction. The effectiveness of the proposed model is extensively validated on two datasets, UK Biobank and ADNI, with a total of 35,388 samples which is one of the largest sample sizes for brain age estimation. It achieves a mean absolute error (MAE) of 2.12 years in UK Biobank and 2.81 years in ADNI, surpassing the performance of existing state-of-the-art approaches. In addition, our method proves valuable in identifying slices that significantly contribute to brain age estimation, facilitating the localization of the age-related regions. In contrast to prior studies, the main contributions of this work can be summarised as follows:
@@
    We propose a novel end-to-end multi-instance learning framework, called DGA-DMIL, which can fully exploit intra- and inter-instance relationships for accurate brain age estimation.
@@
    To address the challenge of heterogeneous aging, we introduce a dual graph attention aggregator to weigh the contribution of different instances of the aging process and to aggregate instance features. In addition, we use contribution scores to visualize the impact of different regions on aging.
@@
    To ameliorate the influence of redundant information, we propose a disentanglement branch that separates age-related and age-independent features to prevent the adverse effect of redundant information on age prediction.
@@ 2.1 Brain age estimation
With superior feature representation abilities, deep learning-based techniques have been successfully applied to medical image analysis in various medical fields including brain age estimation [9, 14, 3, 13, 22, 29, 24]. Most existing brain age estimation methods rely on CNN or Transformer architectures to extract features from 2D MRI slices or 3D volumes. One notable approach adopts the two-stage age network (TSAN) to estimate a rough age at the first stage, followed by a second stage that refines the estimation for improved accuracy [8]. This approach achieves an MAE of 2.43 years on a dataset of 6,586 samples. Another lightweight network, SFCN, proposes an architecture without fully connected layers in brain estimation [41] and achieves an MAE of 2.14 years on a larger dataset of 14,503 samples. In addition, a global-local transformer model is introduced to capture both global contexts from the entire slices and local fine-grained features from specific patches [21]. By using attention mechanisms to combine these features, it achieves an MAE of 2.70 years on a dataset of 8,379 samples.
To the best of our knowledge, our approach is the first attempt to use the MIL method for age prediction based on brain sMRI data. Notably, in this study, our proposed method allows us to exploit the attention scores within and between multiple brain slices to estimate brain age. In addition, we introduce a decoupling branch to effectively separate age-related features from age-independent features, further improving the performance of age prediction. Our method achieved an MAE of 2.12 years on the UK Biobank dataset of 35,291 individuals, outperforming existing approaches.
@@ 2.2 Multi-instance learning
Multi-instance learning (MIL) is a type of weakly supervised learning in which labels are only assigned to a bag of instances, while the labels of instances within the bag remain unknown [50]. MIL can accurately predict the label of the bag by taking into account the effect of the instance on the final result and provide insights on the contributions of intra- and inter-instance [18, 10]. Thus, it has been widely applied in various domains, including medical imaging [42]. Ilse et al. propose an attention-based MIL pooling mechanism focusing on cancer detection in histopathological images [26]. Chikontwe combined MIL with contrastive learning to identify discriminative features for rapid diagnosis of COVID-19 [10]. Marx et al. apply MIL to predict age based on the hippocampal slices [40]. However, it is limited to the hippocampal region, neglecting important information in other brain areas.
@@ 2.3 Decoupled Representation Learning
Disentangled representation learning aims to capture the explanatory factors from diverse data variations, and has attracted considerable attention in recent years. It has shown great promise in the field of computer vision [32, 23]. Previous studies exploit annotated data to separate representations into task-relevant and task-irrelevant features [23]. One notable approach, InfoGAN [7], maximizes the mutual information between hidden and data variables to successfully disentangle latent factors and thus improve performance. Recently, Hu et al. [25] introduce the adversarial autoencoder framework to decouple shared and complementary features across different modalities for accurate brain age prediction. There are notable differences in our proposed method with [25]. (1) Architecture: We adopt an end-to-end non-generative network framework. (2) Decoupling goal: Our research focuses on decoupling age-related and age-independent features, which is particularly beneficial for brain age estimation.
@@ 3.1.1 Problem Formulation
In traditional classification or regression tasks, the goal is to build a model that can accurately predict the target variable, denoted as $y\in R$, given an instance represented by $x\in R^{D}$ [26]. However, in the context of MIL, the input to the model is a collection of instances called a bag, which is represented as $X_{i}=\left\{x_{i,1},x_{i,2},...,x_{i,k}\right\}\in R^{K*D}$, where K is the number of instances within the bag, and D represents the dimension of each instance. There exists a label $y\in R$ at the bag level. The main goal is to gain insight into the inherent properties and contributions of individual instances within a collection of instances for the overall label assigned to the collection. Ultimately, the model aims to make predictions about the label $y$ for the given collection of instances. The formulation of MIL for a regression task is expressed as,
@@ 3.2 CNN backbone
To analyze sMRI data, we adopt a CNN network as the underlying architecture to extract initial features from instances, as shown in Figure 2. To prevent potential overfitting issues associated with deep networks, we employ the VGG13[43] model as our CNN backbone, which is a well-established architecture widely used in medical imaging. The VGG13 model consists of eight blocks, each comprising a $3\times 3$ convolutional layer with 1 padding, a batch normalization layer, and a ReLU activation layer. After every two blocks, a maximum pooling layer with a $2\times 2$ kernel size and a stride of 2 is applied to gradually reduce the size of the spatial feature map. The number of channels utilized in each block is $[64,128,256,512]$.
@@ 3.2.2 Dual Graph Attention Aggregator
Inspired by the attention-based MIL pooling method developed in [26], we propose a novel MIL module, the dual graph attention aggregator (as illustrated in Figure 4), building on the graph attention mechanism. The dual graph attention aggregator consists of a spacial-level feature aggregation network $A_{\theta,S}(R^{K*C*H*W}\to R^{K*C})$ and an instance-level feature aggregation network $A_{\theta,I}(R^{K*C}\to R^{C})$. In the aggregator $A_{\theta,S}$, our goal is to aggregate spatial features from the feature map $e_{i,j}$ into channel features. By passing the feature map $e_{i,j}$ through the GAT block, which consists of GAT, LN, and full connection (FC) layers to reduce the channel dimension to 1, and a softmax layer for normalization, we obtain the final spatial score $g_{i,j}$. To compute the spatial score $g_{i,j}\in R^{1*H*W}$, we perform an element-wise multiplicative convolution of each branch, taking the output layers of the previous layer as input. Mathematically, $g_{i,j}^{{}^{\prime}}=g_{i,j}*e_{i,j}\in R^{C}$, which we denote as $g_{i,j}$ for consistency. In particular, our approach differs from simple global average pooling of the initial features as we adopt the graph attention mechanism to assign different weights within the spatial feature map.